a druggist has an 18% solution of argyol. How much of this solution and how much water must be mixed together to make 10 liters of a 12% solution?
Answer by stanbon(75887)

a druggist has an 18% solution of argyol. How much of this solution and how much water must be mixed together to make 10 liters of a 12% solution?
---------
Equation:
water + water = water
0.82x + 1*(10-x) = 0.88*10
------
Multiply thru by 100 to get:
82x + 100*10 - 100x = 88*10
-18x = -12*10
x = 6 2/3 liters (amt of water needed)
==========================================
Cheers,
Stan H.
---------------------